1. Hospital records show that of patients suffering

from a chronic disease 75% die of it.

What is the probability that of 6 randomly

selected patient 4 will recover.

2. A die is tossed 3 times. What is the probability of:

a. One five turning up

b. 3 fives turning up

c. No five turning up.


1
Expert's answer
2021-01-26T04:05:35-0500

1. P(die) = 75 % = 0.75

P(recover) = 1 - 0.75 = 0.25

"P(4 \;will \;recover) = C^n_xp^xq^{n-x} \\

= C^6_4(0.25)^4(0.75)^2 \\

= 0.0329"

2. a. x =1

"P(X=1) = C^n_xp^xq^{n-x} \\

=C^3_0(\frac{1}{6})^1(\frac{5}{6})^2 \\

= \frac{75}{216} \\

= 0.34722"

b. x=3

"P(X=3) = C^n_xp^xq^{n-x} \\

=C^3_3(\frac{1}{6})^3(\frac{5}{6})^0 \\

= \frac{1}{216} \\

= 4.6296 \times 10^{-3}"

c. x= 0

"P(X=0) = C^n_xp^xq^{n-x} \\

= C^3_0(\frac{1}{6})^0(\frac{5}{6})^3 \\

= \frac{125}{216} \\

= 0.5787"